The mean compression of each spring is 3.5 × 10–3 m
Calculate the spring constant of each spring in the mattress.
Give the unit.
1
Expert's answer
2020-12-17T07:25:52-0500

The mean force on each spring in the mattress is 0.49 N


"k=\frac{F}{x}\\k=\frac{0.49}{3.5\cdot10^{-3}}=140\frac{N}{m}"